OpenStudy (anonymous):

A single tree produces about 2.6 × 102 lb of oxygen each year. The Amazon rainforest has about 3.9 × 1011 trees. About how many pounds of oxygen are produced in the rainforest each day ? 2.8 × 10 lb

OpenStudy (anonymous):

the answer is an exponent

OpenStudy (anonymous):

So to multiply standard form we multiply the 2 numbers and the 2 indices separately. 2.6*3.9=10.14 x^2 * x^11 = x^13 So we've got 10.14*10^13 However Standard form must be 1≤x≤10 therefore we need to divide that by 10 but to get the same number we add one onto the power of 10. So the answer is: 1.014*10^14 lbs of oxygen are produced in the rainforest each year.

OpenStudy (anonymous):

but for the answer I just need to answer in an exponent